Solve 4^2t = 5^4t+4]
I get -2.5803 but i think I did something wrong.

Found 2 solutions by stanbon, solver91311:
Answer by stanbon(57203)

Solve 4^2t = 5^4t+4]
---------
16^t = (5^4]^(t+1)
----
t*log(16) = (t+1)log(625)
----
t(log(16)) = t*log(625) + log(625)
---
t[log(16)-log(625)] = log(625)
---
t = -1.7565
====================
Cheers,
Stan H.
====================

Answer by solver91311(16868)



I have to presume you mean



Because the only way I know to solve



Is to find an approximate root by numerical methods.

Be that as it may, your intuition is spot on.



(order of magnitude difference)



Take the log of both sides (any base, doesn't matter)



Use

to write:



Multiply both sides by



Just for neatness sake, let

Then



The rest is just calculator (or better, spreadsheet) work.

John

My calculator said it, I believe it, that settles it
The Out Campaign: Scarlet Letter of Atheism